A neutralization experiment using NaOH and HCl produced 986 J. The limiting reactant was NaOH, which was present in 1.94×10-2 moles.

Calculate the enthalpy change of neutralization (in kJ/mol).


Give your answer to three significant figures.


1
Expert's answer
2020-08-25T12:19:39-0400

"NaOH+HCl \to NaCl +H_2O"

This is the neutralisation reaction.

So, the enthalply change to be calculated will be on enthalpy change per mole of "NaOH" reacted( As it is limiting reagent)

Net Enthalply change given "(H)=986 \ J"

Moles reacted "(N) =1.94\times 10^{-2}"

Heat of reaction "(H_r) =\frac{H}{N}=\frac{986}{1.94\times 10^{-2}}J=50824.74J=50.824\ KJ""=50.8 \ KJ" (upto 3 significant figures)